a stretched sonometer wire resonates at a frequency of 350 hz and at the next higher frequency of 420 hz .Fundamental frequency of this wire is

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
(n+1)/n = 420/350=6/5
1/n+1 = 6/5 
1/n = 1/5
n=5
350/5 = 70 Hz is findamental frequency , 
Hope thia hwlps 
Good Luck